# Positive Semidefinite Matrix Supermartingales

### v4

@@ 5.1 Tightness of Matrix Markov’s and Ville’s Inequalities
[…]
@@
    If $\mathbf{A}\in\mathcal{A}_{\alpha}$, then $\mathcal{C}(\mathbf{A})\subseteq\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$ if and only if $\alpha^{\prime}\leqslant d(\mathrm{tr}\mathbf{A})^{-1}\in[\frac{\alpha}{d},\frac{\alpha}{\operatorname{r}(\mathbf{A}^{-1})}]$;
@@
Let us briefly unpack these statements above. The first two state that (MMI) and trace-(MI) are mutually incomparable unless one takes the threshold matrix $\mathbf{A}$ to be a multiple of $\mathbf{I}$, in which case (MMI) is weaker than trace-(MI). Statements 3 and 4 indicate that (MMI) and trace-(MI) can simulate each other only by increasing the error probability $\alpha$, and that if $\mathbf{A}^{-1}$ is chosen to have a low effective rank, (MMI) can easily simulate trace-(MI) whereas trace-(MI) pays a price up to $d$ to simulate (MMI). Statements 5 and 6 imply respectively that, while one can always simulate quadratic form-(MI) with (MMI) by infinitesimally increasing the error probability $\alpha$, the converse is never true. Proposition 5.1 is proved in Section B.4.
@@ 5.2 Stepwise Scalarization
The proof of this lemma can be found in Section B.5. We have numerous remarks regarding this lemma.
@@
Second, the divergence from the common assumption (67) to the matrix-valued (68) and scalar-valued (69) test supermartingale indicates that for some testing problems, two different routes can be pursued: our “matrix-native” approach, and the scalarization approach to the test process. Examples in Section 3.3 that fit in the assumption (67) include Examples 3.8 and 3.9 which Howard et al. [2020, Section 6.5] originally prove, but also Example 3.10: To see that Example 3.10 is in the form of (67),
@@
which is exactly what is done by Wang and Ramdas [2024, Theorem 4.2] to derive a sharp closed-form matrix empirical Bernstein inequality. One of the crucial ways the two approaches differ is that with the scalar-valued test process (69), one rejects the null when the process surpasses $d/\alpha$; whereas with the matrix-valued test process (68), one rejects when the process $\npreceq\mathbf{A}$ where $\mathbf{A}$ is a matrix chosen freely beforehand, only subject to the “size constraint” that $\mathrm{tr}(\mathbf{A}^{-1})=\alpha$. Indeed, the extra degrees of freedom that come with $\mathbf{A}$ play a crucial role as we compare the two approaches empirically with the self-normalizing example (Example 3.8) in Section 6.
Third, the assumption (67) for scalarization being more restrictive than our Lemma 3.5, for some testing problems our concept of matrix-valued test supermartingales is unavoidable. These include Examples 3.7 and 3.12, to which the stepwise scalarization Lemma 5.2 is not applicable. Their one-step condition $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}$ leads naturally to a matrix-valued test supermartingale, and scalar-valued test supermartingales might be obtained by taking the trace or quadratic form after the formation of the matrix-valued one, which is not ideal according to Section 5.1.
@@ 5.3 Randomized Ahlswede-Winter Bounds
$\mathbb{P}(X\leqslant Ua)\leqslant(\mathbb{E}X)a^{-1}$ (72)
@@
It is clear that (73) generalizes the scalar (72) and tightens (MMI). The randomization factor $U$, we note, can be any random variable that is stochastically larger than $\operatorname{Unif}_{(0,1)}$ (“super-uniform”), that is, $\mathbb{P}(U\leqslant x)\leqslant x$ for all $x\geqslant 0$. More generally, we shall present and prove a stronger version of Theorem 5.3 in Appendix A, where the scalar randomization factor $U$ is further replaced by a PSD random matrix whose distribution belongs to a class of “super-uniform” measures on $\mathcal{S}_{d}^{+}$. The generalized version of Theorem 5.3 in Appendix A also reveals the equality condition for (MMI), randomized or not.
@@
We omit the proof since it is a straightforward application of Theorem 5.3 to the moment bound Lemma B.2. We next discuss the implication for (MVI) as the randomization here raises some subtleties in its two equivalent forms, the time-uniform one and the stopped one.
@@
First, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$ due to Part 3 of Theorem 3.3. We apply Theorem 5.3 to obtain the first inequality. For the second inequality, we define the stopping time $\nu:=\inf\{n:\mathbf{Y}_{n}\npreceq\mathbf{A}\}$ and apply (75) to the stopping time $\nu\wedge\tau$. ∎
@@ 6 Experiments: To Scalarize or Not to Scalarize
We conduct the comparison with the self-normalized Example 3.8, and consider the two corresponding sequential test algorithms, Matrix and Scalar. More specifically, both receive the stream of i.i.d matrices $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ in $\mathcal{S}_{d}$ with mean $\mathbf{M}$ and variance upper bound $\mathbf{I}$ and sequentially test the null hypothesis $H_{0}:\mathbf{M}=0$. Matrix computes the matrix-valued process
@@
Under the null, both Matrix and Scalar safely control the type 1 error (the probability of falsely rejecting the null) under the predefined level $\alpha$ due to (MVI) and (VI). For power comparison, we shall first discuss the case when only $n=1$ observation is revealed, and then move on to the case with an indefinite sequence of observations, with examples where Matrix outpowers Scalar under decreasingly restrictive conditions and finally no restriction at all.
@@ 6.2 Comparison with a Sequence of Observations
with common deterministic projection matrices $\mathbf{P}_{1},\dots,\mathbf{P}_{d}$. For each $k$, $\lambda_{k1},\lambda_{k2},\dots\overset{\mathrm{iid}}{\sim}\pi_{k}$ with mean $\mathbb{E}\lambda_{k1}=\lambda_{k}$ and variance $\mathbf{Var}\lambda_{k1}=\sigma_{k}^{2}\leqslant 1$. Then, the variance $\mathbf{Var}(\mathbf{X}_{1})=\sigma_{1}^{2}\mathbf{P}_{1}+\dots+\sigma_{d}^{2}\mathbf{P}_{d}\preceq\mathbf{I}$. The matrix-valued process that Matrix computes is
$\mathbf{Y}_{n}^{\textsf{SN}}=\sum_{k=1}^{d}Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})\mathbf{P}_{k},$ (85)
@@
$L_{n}^{\textsf{SN}}=\sum_{k=1}^{d}Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})=\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}}).$ (86)
In both expressions the function $Y_{n}^{\textsf{SN}}(\cdot)$ denotes the self-normalized test process for scalar observations,
$Y_{n}^{\textsf{SN}}(x_{1},\dots,x_{n})=\exp\left(-\frac{1}{3}\sum_{i=1}^{n}\gamma_{i}^{2}+\sum_{i=1}^{n}\gamma_{i}x_{i}-\frac{1}{6}\sum_{i=1}^{n}\gamma_{i}^{2}x_{i}^{2}\right).$ (87)
The scalar component process $Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})$ grows exponentially fast if $\pi_{k}$ has mean $\lambda_{k}\neq 0$, and shrinks to 0 if $\lambda_{k}=0$; moreover, the rate in which $Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})$ grows depends exponentially on the size $|\lambda_{k}|$. Consequently, as long as there is any mean disparity between the component distributions $\pi_{1},\dots,\pi_{d}$, say $|\lambda_{1}|,\dots,|\lambda_{d-1}|<|\lambda_{d}|$, in the long run the process $\mathbf{Y}_{n}^{\textsf{SN}}$ will have a low effective rank with $\operatorname{r}(\mathbf{Y}_{n}^{\mathsf{SN}})={\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}\to 1$, as the component $Y_{n}^{\textsf{SN}}(\lambda_{d1},\dots,\lambda_{dn})\mathbf{P}_{d}$ will exponentially outgrow the rest. In this case if Matrix correctly anticipates the largest component $\mathbf{P}_{d}$, by picking again the threshold matrix
@@
with small $\varepsilon$ it will outperform Scalar: it only takes $Y_{n}^{\textsf{SN}}(\lambda_{d1},\dots,\lambda_{dn})$ to surpass $(\alpha-\varepsilon)^{-1}\approx 1/\alpha$ for Matrix to reject, whereas it takes $L_{n}^{\textsf{SN}}\approx Y_{n}^{\textsf{SN}}(\lambda_{d1},\dots,\lambda_{dn})$ to surpass $d/\alpha$ for Scalar to reject. Similar reasoning applies as long as Matrix knows a subset of the $d$ components with larger means.
@@
smart Matrix $>$ naive Matrix $\approx$ Scalar
@@
smart Matrix $>$ naive Matrix $>$ Scalar.
@@
7 Conclusion
Acknowledgement
Appendix A On the Randomized Matrix Markov Inequality
@@
Theorem A.2 (Generalized Uniformly Randomized Matrix Markov Inequality).
@@
Appendix B Omitted Proofs
B.1 Dominated Integrability Lemma
@@
Lemma B.1 (Dominated Integrability).
@@
B.2 Proofs for Forward Supermartingales
@@
Note that
@@
$\iff$ For all $n$, $\mathbf{Y}_{n-1}-\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\succeq 0$ (100)
$\iff$ For all $n$ and $\mathbf{v}$, $\mathbf{v}^{\mathsf{T}}(\mathbf{Y}_{n-1}-\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1}))\mathbf{v}\geqslant 0$ (101)
$\iff$ For all $n$ and $\mathbf{v}$, $\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n-1}\mathbf{v}-\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}|\mathcal{F}_{n-1})\geqslant 0$ (102)
[…]
@@
This allows us to use Lemma B.1,
@@
=\ = $1+\mathbb{E}\left(\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/3\mid\mathcal{F}_{n-1}\right)\preceq 1+\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\preceq\exp\left(\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\right),$ (107)
@@
Let us replicate the key steps of the proof of Lemma 3(d) in Howard et al. [2020], which is also a self-normalization argument. Due to the symmetry of $\mathbf{X}_{n}-\mathbf{M}_{n}$, we introduce a Rademacher random variable $\epsilon$ independent from $\mathcal{F}_{n}$,
@@
Note that the function $f(x)=\frac{\exp(x-x^{2}/2)+\exp(-x-x^{2}/2)}{2}$ satisfies $f(x)\leqslant 1$ for all $x\in\mathbb{R}$, we see from the transfer rule and Lemma B.1 that
@@
B.3 Proofs for Backward Submartingales
Lemma B.2 (Contraction of the Matrix Variance).

### v6

@@ 5.1 Tightness of Matrix Markov’s and Ville’s Inequalities
Recall that a PSD supermartingale $\{\mathbf{Y}_{n}\}$ can be converted into a scalar nonnegative supermartingale via either taking the trace $\{\mathrm{tr}\mathbf{Y}_{n}\}$ or a quadratic form $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$. We thus discuss the tightness of (MVI) in this subsection in comparison to the scalar (VI) on these transformations. Since (MVI) is proved by applying (MMI) on a stopped random matrix, it suffices to compare (MMI) versus (MI) on a PSD random matrix $\mathbf{X}$. Without loss of generality, we assume that $\mathbb{E}\mathbf{X}=\mathbf{I}$. Denote by $\mathcal{C}_{\mathrm{tr}}(\alpha)$ the concentration region by applying (MI) to $\mathrm{tr}\mathbf{X}$:
@@
    If $\mathbf{A}\in\mathcal{A}_{\alpha}$, then $\mathcal{C}(\mathbf{A})\subseteq\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$ if and only if $\alpha^{\prime}\leqslant d(\mathrm{tr}\mathbf{A})^{-1}$ $\in(0,\frac{\alpha}{\operatorname{r}(\mathbf{A})}]$;
@@
Let us briefly unpack these statements above. The first two state that (MMI) and trace-(MI) are mutually incomparable unless one takes the threshold matrix $\mathbf{A}$ to be a multiple of $\mathbf{I}$, in which case (MMI) is weaker than trace-(MI). Statements 3 and 4 indicate that (MMI) and trace-(MI) can simulate each other only by increasing the error probability $\alpha$ substantially. Statements 5 and 6 imply respectively that, while one can always simulate quadratic form-(MI) with (MMI) by infinitesimally increasing the error probability $\alpha$, the converse is never true. Proposition 5.1 is proved in Section 8.4.
@@ 5.2 Stepwise Scalarization
The proof of this lemma can be found in Section 8.5. We have numerous remarks regarding this lemma.
@@
Second, the divergence from the common assumption (66) to the matrix-valued (67) and scalar-valued (68) test supermartingale indicates that for some testing problems, two different routes can be pursued: our “matrix-native” approach, and the scalarization approach to the test process. Examples in Section 3.3 that fit in the assumption (66) include Examples 3.9 and 3.10 which Howard et al. [2020, Section 6.5] originally prove, but also Example 3.11: To see that Example 3.11 is in the form of (66),
@@
which is exactly what is done by Wang and Ramdas [2024, Theorem 4.2] to derive a sharp closed-form matrix empirical Bernstein inequality. One of the crucial ways the two approaches differ is that with the scalar-valued test process (68), one rejects the null when the process surpasses $d/\alpha$; whereas with the matrix-valued test process (67), one rejects when the process $\npreceq\mathbf{A}$ where $\mathbf{A}$ is a matrix chosen freely beforehand, only subject to the “size constraint” that $\mathrm{tr}(\mathbf{A}^{-1})=\alpha$. Indeed, the extra degrees of freedom that come with $\mathbf{A}$ play a crucial role as we compare the two approaches empirically with the finite-variance self-normalizing example (Example 3.9) in Section 6.
Third, the assumption (66) for scalarization being more restrictive than our Lemma 3.5, for some testing problems our concept of matrix-valued test supermartingales is unavoidable. These include Example 3.7 and 3.13, to which the stepwise scalarization Lemma 5.2 is not applicable. Their one-step condition $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}$ leads naturally to a matrix-valued test supermartingale, and scalar-valued test supermartingales might be obtained by taking the trace or quadratic form after the formation of the matrix-valued one, which is not ideal according to Section 5.1.
@@ 5.3 Randomized Ahlswede-Winter Bounds
$\mathbb{P}(X\geqslant Ua)\leqslant(\mathbb{E}X)a^{-1}$ (71)
@@
It is clear that (72) generalizes the scalar (71) and tightens (MMI). The randomization factor $U$, we note, can be any random variable that is stochastically larger than $\operatorname{Unif}_{(0,1)}$ (“super-uniform”), that is, $\mathbb{P}(U\leqslant x)\leqslant x$ for all $x\geqslant 0$. More generally, we shall present and prove a stronger version of Theorem 5.3 in Section 7, where the scalar randomization factor $U$ is further replaced by a PSD random matrix whose distribution belongs to a class of “super-uniform” measures on $\mathcal{S}_{d}^{+}$. The generalized version of Theorem 5.3 in Section 7 also reveals the equality condition for (MMI), randomized or not.
@@
We omit the proof since it is a straightforward application of Theorem 5.3 to the moment bound Lemma 8.2. We next discuss the implication for (MVI) as the randomization here raises some subtleties in its two equivalent forms, the time-uniform one and the stopped one.
@@
First, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$ due to Part 3 of Theorem $3.$ We apply Theorem 5.3 to obtain the first inequality. For the second inequality, we define the stopping time $\nu:=\inf\{n:\mathbf{Y}_{n}\npreceq\mathbf{A}\}$ and apply (74) to the stopping time $\nu\wedge\tau$. ∎
@@ 6 Experiments: To Scalarize or Not to Scalarize
We conduct the comparison with the finite-variance self-normalized Example 3.9, and consider the two corresponding sequential test algorithms, Matrix and Scalar. More specifically, both receive the stream of i.i.d matrices $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ in $\mathcal{S}_{d}$ with mean $\mathbf{M}$ and variance upper bound $\mathbf{I}$ and sequentially test the null hypothesis $H_{0}:\mathbf{M}=0$. Matrix computes the matrix-valued process
@@
Under the null, both Matrix and Scalar safely control the type 1 error (the probability of falsely rejecting the null) under the predefined level $\alpha$ due to (MVI) and (VI). For power comparison, we shall first discuss the case when only $n=1$ observation is revealed, and then move on to the case with an indefinite sequence of observations, with examples where Matrix outpowers Scalar under decreasingly restrictive conditions and finally no restriction at all. We shall specify the choice of the $\{\gamma_{n}\}$ sequence as needed.
@@ 6.2 Comparison with a Sequence of Observations
with common deterministic projection matrices $\mathbf{P}_{1},\dots,\mathbf{P}_{d}$. For each $k$, $\lambda_{k1},\lambda_{k2},\dots\overset{\mathrm{iid}}{\sim}\pi_{k}$ with mean $\mathbb{E}\lambda_{k1}=\lambda_{k}\geqslant 0$ and variance $\mathbf{Var}\lambda_{k1}=\sigma_{k}^{2}\leqslant 1$. Then, the variance $\mathbf{Var}(\mathbf{X}_{1})=\sigma_{1}^{2}\mathbf{P}_{1}+\dots+\sigma_{d}^{2}\mathbf{P}_{d}\preceq\mathbf{I}$.
We fix a deterministic sequence of positive scalars $\gamma_{n}=n^{-1/2}$. For each $k$, we define the auxiliary “component process” $\{F_{kn}\}_{n\geqslant 0}$ as the scalar finite-variance self-normalized test process with $\gamma_{n}=n^{-1/2}$ as the weight sequence when observing data $\lambda_{k1},\lambda_{k2},\dots$:
$F_{kn}=\exp\left(-\frac{1}{3}\sum_{i=1}^{n}\gamma_{i}^{2}+\sum_{i=1}^{n}\gamma_{i}\lambda_{ki}-\frac{1}{6}\sum_{i=1}^{n}\gamma_{i}^{2}\lambda_{ki}^{2}\right).$ (84)
The matrix-valued process that Matrix computes can then be written as
$\mathbf{Y}_{n}^{\textsf{SN}}=\sum_{k=1}^{d}F_{kn}\mathbf{P}_{k},$ (85)
@@
$L_{n}^{\textsf{SN}}=\sum_{k=1}^{d}F_{kn}=\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}}).$ (86)
Note that $\{F_{kn}\}$ satisfies
$F_{kn}=F_{k(n-1)}\cdot\exp\left(\frac{\lambda_{kn}}{\sqrt{n}}-\frac{2+\lambda_{kn}^{2}}{6n}\right).$ (87)
For large $n$, the first $n^{-1/2}$ term inside the $\exp(\cdot)$ above dominates the second $n^{-1}$ term. We see that each scalar component process $\{F_{kn}\}$ eventually grows exponentially fast if $\lambda_{k}=\mathbb{E}\lambda_{kn}>0$, and shrinks to 0 if $\lambda_{k}=0$; moreover, the exponential growth rate of $\{F_{kn}\}$ depends polynomially on the effect size $\lambda_{k}$. Consequently, as long as there is any mean disparity between the component distributions $\pi_{1},\dots,\pi_{d}$, say $\lambda_{1},\dots,\lambda_{d-1}<\lambda_{d}$, in the long run the process $\mathbf{Y}_{n}^{\textsf{SN}}$ will have a low effective rank with $\operatorname{r}(\mathbf{Y}_{n}^{\mathsf{SN}})={\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}\to 1$, as the component $F_{dn}\mathbf{P}_{d}$ will exponentially outgrow the rest $F_{1n}\mathbf{P}_{1},\dots,F_{(d-1)n}\mathbf{P}_{d-1}$. In this case if Matrix correctly anticipates the largest component $\mathbf{P}_{d}$, by picking again the threshold matrix
@@
with small $\varepsilon$ it will outperform Scalar: it only takes $F_{dn}$ to surpass $(\alpha-\varepsilon)^{-1}\approx 1/\alpha$ for Matrix to reject, whereas it takes $L_{n}^{\textsf{SN}}\approx F_{dn}$ to surpass $d/\alpha$ for Scalar to reject. Similar reasoning applies as long as Matrix knows a subset of the $d$ components with larger means. We finally remark that we can lift the assumption $\lambda_{k}\geqslant 0$ here, by using the “hedging” argument outlined in Remark 3.8. Let $\mathbf{Y}_{n}^{+}$ and $\mathbf{Y}_{n}^{-}$ be the processes Matrix would use with $\gamma_{n}=n^{-1/2}$ and $\gamma_{n}=-n^{-1/2}$ respectively. Then, if $|\lambda_{1}|,\dots,|\lambda_{d-1}|<|\lambda_{d}|$, the spectrum of $(\mathbf{Y}_{n}^{+}+\mathbf{Y}_{n}^{-})/2$ will concentrate on the $\mathbf{P}_{d}$ direction, too.
@@
smart Matrix >> naive Matrix $\approx$ Scalar
@@
smart Matrix >> naive Matrix >> Scalar.
@@
7 On the Randomized Matrix Markov Inequality
@@
Theorem 7.2 (Generalized Uniformly Randomized Matrix Markov Inequality).
@@
8 Omitted Proofs
8.1 Dominated Integrability Lemma
@@
Lemma 8.1 (Dominated Integrability).
@@
8.2 Proofs for Forward Supermartingales
@@
(1) Note that $\mathbf{A}\succeq 0$ if and only if $\mathbf{v}^{\mathsf{T}}\mathbf{A}\mathbf{v}\geqslant 0$ for all $\mathbf{v}\in\mathbb{Q}^{d}$ due to the continuity of the quadratic form. We have,
@@
$\iff$ $\mathbb{P}\left(\bigcap_{n}\{\mathbf{Y}_{n-1}-\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\succeq 0\}\right)=1$ (100)
$\iff$ $\mathbb{P}\left(\bigcap_{n}\bigcap_{v\in\mathbb{Q}^{d}}\{\mathbf{v}^{\mathsf{T}}(\mathbf{Y}_{n-1}-\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1}))\mathbf{v}\geqslant 0\}\right)=1$ (101)
$\iff$ $\mathbb{P}\left(\bigcap_{n}\bigcap_{v\in\mathbb{Q}^{d}}\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n-1}\mathbf{v}-\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}|\mathcal{F}_{n-1})\geqslant 0\}\right)=1$ (102)
$\iff$ For all $\mathbf{v}\in\mathbb{Q}^{d}$, $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale (103)
$\iff$ For all $\mathbf{v}\in\mathbb{R}^{d}$, $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale. (104)
(2) Since $\mathbf{A}_{n}=\sum_{i=1}^{n}(\mathbb{E}(\mathbf{Y}_{i}|\mathcal{F}_{i-1})-\mathbf{Y}_{i-1})$,
$\{\mathbf{Y}_{n}\}$ is a supermartingale (105)
$\iff$ For all $n$, $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})-\mathbf{Y}_{n-1}\preceq 0$ (106)
$\iff$ $\text{For all $n$, $\mathbf{A}_{n}-\mathbf{A}_{n-1}\preceq 0$.}\qed$ (107)
@@
This allows us to use Lemma 8.1,
@@
$=\$$1+\mathbb{E}\left(\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/3\mid\mathcal{F}_{n-1}\right)\preceq 1+\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\preceq\exp\left(\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\right),$ (111)
@@
Let us replicate the key steps of the proof of Lemma 3(d) in Howard et al. [2020], another self-normalization-style bound derived this time by the classical symmetrization argument. Due to the symmetry of $\mathbf{X}_{n}-\mathbf{M}_{n}$, we introduce a Rademacher random variable $\epsilon$ independent from $\mathcal{F}_{n}$,
@@
Note that the function $f(x)=\frac{\exp(x-x^{2}/2)+\exp(-x-x^{2}/2)}{2}$ satisfies $f(x)\leqslant 1$ for all $x\in\mathbb{R}$, we see from the transfer rule and Lemma 8.1 that
@@
8.3 Proofs for Backward Submartingales
Lemma 8.2 (Contraction of the Matrix Variance).